# Precompactness in matrix weighted Bourgain-Morrey spaces

### v1

@@ 4 Dyadic average operator conditions for precompact sets
$\approx W(Q_{j,k})^{1/t-1/p}2^{-jn/p}2^{(j+a)\tilde{d}/p}\left(\frac{1}{|(j+a)_{\operatorname{pa}}Q_{j,k}|}\right.$
×$\quad\times\left.\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\frac{1}{|(j+a)_{\operatorname{pa}}Q_{j,k}|}\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\vec{f}(y){\mathrm{d}}y\right|^{p}{\mathrm{d}}x\right)^{1/p}.$
Then by Lemma 4.4 and $\tilde{d}/p-n/p-\beta(1/t-1/p)\leq 0$, we obtain
@@
$\lesssim W(Q_{j,k})^{1/t-1/p}2^{-jn/p}2^{(j+a)\tilde{d}/p}$
×$\quad\times\left(\frac{1}{|(j+a)_{\operatorname{pa}}Q_{j,k}|}\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\vec{f}(x)\right|^{p}{\mathrm{d}}x\right)^{1/p}$
@@
$\lesssim 2^{(j+a)(\tilde{d}/p-n/p)}2^{-(j+a)\beta(1/t-1/p)}W((j+a)_{\operatorname{pa}}Q_{j,k})^{1/t-1/p}$
×$\quad\times\left(\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\vec{f}(x)\right|^{p}{\mathrm{d}}x\right)^{1/p}$
$\lesssim W((j+a)_{\operatorname{pa}}Q_{j,k})^{1/t-1/p}\left(\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\vec{f}(x)\right|^{p}{\mathrm{d}}x\right)^{1/p}$
where in the third step we use the following fact. Since $W((j+a)_{\operatorname{pa}}Q)\leq 2^{(j+a)\beta}W(Q)$ and $1/t-1/p\leq 0$, we have
$W(Q)^{1/t-1/p}\leq 2^{-(j+a)\beta(1/t-1/p)}W((j+a)_{\operatorname{pa}}Q)^{1/t-1/p}.$
@@
Theorem 4.6
Let $1\leq p<t<r<\infty$ Let $W\in\mathcal{A}_{p}$ has the $\mathcal{A}_{p}$-dimension $\tilde{d}\in[0,n)$ with the doubling exponent $\beta$. Let $(\tilde{d}-n+\beta)r/p+n-\beta r/t<0$. A subset $\mathcal{F}$ of $M_{p}^{t,r}(W)$ is totally bounded if and only if the following conditions hold:
@@
where $E_{d,a}$ is same as (4.2).
  • Proof
    The sufficiency is due to Theorem 4.1. Now we prove the necessity. Assume that $\mathcal{F}$ of $M_{p}^{t,r}(W)$ is totally bounded. For any given $\varepsilon>0$, there exists $\{\vec{f}_{k}\}_{k=1}^{N_{0}}\subset\mathcal{F}$ such that $\{\vec{f}_{k}\}_{k=1}^{N_{0}}$ is an $\varepsilon$-net of $\mathcal{F}$, that is, for any $\vec{f}\in\mathcal{F}$, there exists $\vec{f}_{k},k\in\{1,2,\ldots,N_{0}\}$ such that $\|\vec{f}-\vec{f}_{k}\|_{M_{p}^{t,r}(W)}<\varepsilon$.
@@
    For any $\varepsilon>0$ and $k\in\{1,2,\ldots,N_{0}\}$, since $r<\infty$, there exists $R_{k}>0$ such that
    $\|\vec{f}_{k}\chi_{B^{c}(0,R_{k})}\|_{M_{p}^{t,r}(W)}<\varepsilon.$
    Taking $R=\max_{k\in{1,\ldots,N_{0}}}\{R_{k}\}$, we have $\|\vec{f}_{k}\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}<\varepsilon.$ Then for any $\vec{f}\in\mathcal{F}$,
    $\|\vec{f}\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}\leq\|(\vec{f}-\vec{f}_{k})\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}+\|\vec{f}_{k}\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}<2\varepsilon.$
@@
$\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}$
$\leq\|E_{d,a}\vec{f}-E_{d,a}\vec{f}_{k}\|_{M_{p}^{t,r}(W)}+\|E_{d,a}\vec{f}_{k}-\vec{f}_{k}\|_{M_{p}^{t,r}(W)}+\|\vec{f}_{k}-\vec{f}\|_{M_{p}^{t,r}(W)}$
$\leq c\|\vec{f}-\vec{f}_{k}\|_{M_{p}^{t,r}(W)}+\varepsilon+\varepsilon$
$\lesssim\varepsilon.$
Remark 4.7
@@
Let $f_{j}=\chi_{B(0,2^{j})}f$ for $j\in\mathbb{N}$. Note that $f$ and $f_{j}$ are radial and symmetric functions. Let $x=(x_{1},0,\ldots,0)$ where $x_{1}=2^{j}+2\times 2^{j}$. Let $s=x_{1}/10$ Then $B(x,s)\subset B(0,2^{j})^{c}$. Then
$\|f-f_{j}\|_{M^{t,\infty}_{p}}$
$\geq cs^{-n(1/t-1/p)}\left(\int_{B(x,s)}|y|^{-np/t}{\mathrm{d}}y\right)^{1/p}$
$\geq cs^{-n(1/t-1/p)}\left(\int_{|x|-|s|}^{|x|+|s|}\eta^{-np/t}\eta^{n-1}{\mathrm{d}}\eta\right)^{1/p}$
@@
This shows that the totally bounded set =$\mathcal{F}=\{f\}$ of $M^{t,\infty}_{p}$ is not uniformly vanishes at infinity.
  • Acknowledgement(s).  The corresponding author Jingshi Xu is supported by the National Natural Science Foundation of China (Grant No. 12161022) and the Science and Technology Project of Guangxi (Guike AD23023002).

### v2

@@ 4 Dyadic average operator conditions for precompact sets
$\approx W(Q_{j,k})^{1/t-1/p}2^{-jn/p}2^{(j+a)\tilde{d}/p}\left(\frac{1}{|(j+a)_{\operatorname{pa}}Q_{j,k}|}\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\frac{1}{|(j+a)_{\operatorname{pa}}Q_{j,k}|}\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\vec{f}(y){\mathrm{d}}y\right|^{p}{\mathrm{d}}x\right)^{1/p}.$
Then by Lemma 4.4 and $\tilde{d}/p-n/p-\tilde{\beta}(1/t-1/p)\leq 0$, we obtain
@@
$\lesssim W(Q_{j,k})^{1/t-1/p}2^{-jn/p}2^{(j+a)\tilde{d}/p}\left(\frac{1}{|(j+a)_{\operatorname{pa}}Q_{j,k}|}\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\vec{f}(x)\right|^{p}{\mathrm{d}}x\right)^{1/p}$
@@
$\lesssim 2^{(j+a)(\tilde{d}/p-n/p)}2^{-(j+a)\tilde{\beta}(1/t-1/p)}W((j+a)_{\operatorname{pa}}Q_{j,k})^{1/t-1/p}\left(\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\vec{f}(x)\right|^{p}{\mathrm{d}}x\right)^{1/p}$
$\lesssim W((j+a)_{\operatorname{pa}}Q_{j,k})^{1/t-1/p}\left(\int_{(j+a)_{\operatorname{pa}}Q_{j,k}}\left|W^{1/p}(x)\vec{f}(x)\right|^{p}{\mathrm{d}}x\right)^{1/p}.$
@@
Theorem 4.7.
Let $1\leq p<t<r<\infty$ . Let $W\in\mathcal{A}_{p}$ with the $\mathcal{A}_{p}$-dimension $\tilde{d}\in[0,n)$. Let $\tilde{\beta}$ be the same with (24). Let $nr/p+\tilde{d}r/p+n-\tilde{\beta}r(1/t-1/p)<0$. A subset $\mathcal{F}$ of $M_{p}^{t,r}(W)$ is totally bounded if and only if the following conditions hold:
@@
where $E_{d,a}$ is same as (19).
Proof.
The sufficiency is due to Theorem 4.1. Now we prove the necessity. Assume that $\mathcal{F}$ of $M_{p}^{t,r}(W)$ is totally bounded. For any given $\epsilon>0$, there exists $\{\vec{f}_{k}\}_{k=1}^{N_{0}}\subset\mathcal{F}$ such that $\{\vec{f}_{k}\}_{k=1}^{N_{0}}$ is an $\epsilon$-net of $\mathcal{F}$, that is, for any $\vec{f}\in\mathcal{F}$, there exists $\vec{f}_{k},k\in\{1,2,\ldots,N_{0}\}$ such that $\|\vec{f}-\vec{f}_{k}\|_{M_{p}^{t,r}(W)}<\epsilon$.
@@
For any $\epsilon>0$ and $k\in\{1,2,\ldots,N_{0}\}$, since $r<\infty$, there exists $R_{k}>0$ such that
$\|\vec{f}_{k}\chi_{B^{c}(0,R_{k})}\|_{M_{p}^{t,r}(W)}<\epsilon.$
Taking $R=\max_{k\in\{1,\ldots,N_{0}\}}\{R_{k}\}$, we have $\|\vec{f}_{k}\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}<\epsilon.$ Then for any $\vec{f}\in\mathcal{F}$,
$\|\vec{f}\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}\leq\|(\vec{f}-\vec{f}_{k})\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}+\|\vec{f}_{k}\chi_{B^{c}(0,R)}\|_{M_{p}^{t,r}(W)}<2\epsilon.$
@@
$\|\vec{f}-E_{d,a}\vec{f}\|_{M_{p}^{t,r}(W)}$ $\leq\|E_{d,a}\vec{f}-E_{d,a}\vec{f}_{k}\|_{M_{p}^{t,r}(W)}+\|E_{d,a}\vec{f}_{k}-\vec{f}_{k}\|_{M_{p}^{t,r}(W)}+\|\vec{f}_{k}-\vec{f}\|_{M_{p}^{t,r}(W)}$
$\leq c\|\vec{f}-\vec{f}_{k}\|_{M_{p}^{t,r}(W)}+\epsilon+\epsilon\lesssim\epsilon.$
Then the proof is complete. ∎
Remark 4.8.
@@
Let $f_{j}=\chi_{B(0,2^{j})}f$ for $j\in\mathbb{N}$. Note that $f$ and $f_{j}$ are radial and symmetric functions. Let $x=(x_{1},0,\ldots,0)$ where $x_{1}=2^{j}+2\times 2^{j}$. Let $s=x_{1}/10$ . Then $B(x,s)\subset B(0,2^{j})^{c}$. Then
$\|f-f_{j}\|_{M^{t,\infty}_{p}}$ $\geq cs^{-n(1/t-1/p)}\left(\int_{B(x,s)}|y|^{-np/t}{\mathrm{d}}y\right)^{1/p}\geq cs^{-n(1/t-1/p)}\left(\int_{|x|-|s|}^{|x|+|s|}\eta^{-np/t}\eta^{n-1}{\mathrm{d}}\eta\right)^{1/p}$
@@
This shows that the totally bounded set $\mathcal{F}=\{f\}$ of $M^{t,\infty}_{p}$ is not uniformly vanishes at infinity.